This 3 part question has been puzzling me
What value of m gives a system
y=1x/2x-1
y=mx-1 with
a)one solution b)an infinitely many solution c) no solution

Answer by Mathtut(3670)

At first I thought A. B and C were choices of answers.........
:
they want you to find m such that it satisfys the following conditions. The problem was written very poorly.
:
A one solution
B.infinite solutions
C no solutions
:
:
A) m could be almost anything except 1^2.......so let m=-1 then we have y=1^2x-1 and y=-1x-1 and those lines will cross in one place as they are perpendicular to each other
B.) in order for this system to have infinite solutions it has to be the same line so m=1^2
C) There is no value of m that will satisfy this condition because the slope must be the same(parallel lines) to have no solutions. The problem is the y intercept would also have to change to be a parallel line because if m is equal to 1^2 we would have the same answer as B( same line-infinite